by TEI TEI

10/14/05 – TTC has inquired about status of project. If the new door is being handled within, then this project is complete. Diana and Mark please advise as to the status of this project.

05/12/05 – We all discussed alternate schemes for a new door to meet state and local requirements and TTC sent this color scheme with 2 options to GBBN for review and for GBBN to forward to the County for approval. The County shall make the final decision as to which way they want to proceed. PDF

04/26/05 – TE recommended provide a new door in the space to meet code requirements from state and city per this PDF. TE must wait for County to approve a way to go on this issue before project is complete.

02/2/05 – TTC talked to Mike Bartlett and found that complaints have subsided since Greg greatly reduced the total amount of air being fed into the space. This makes sense since the space was being overcooled because the thermostat was on the second floor in the courtroom. Mike and I agree that this should work until warmer weather where the airflow into the space will have to be increased and further substantiates the need for a revised VAV system in this space. As it right now, Mike and Greg are acting as a manual VAV system by increasing and decreasing air into the space as needed. This also shows us that if you are really crunched on money, you may be able to only put a couple of VAV boxes into the first floor area and begin to achieve some control with the hopes that a full redesign can be done later.

02/25/05 – Client reported that space is still not conditioned during the evening time properly. The daytime loads can still never be fixed without a new VAV system and HW reheat (see SOSS TE-115)

02/10/05 – TE has not seen a balance report or payapp for the additional work.

02/01/05 – MB reported today that Room 169 reported no comfort problems with the space except when the main ahu turned on. PHB still has to balance the new air device. TE to punchlist this job shortly.

01/31/05 – MB reported that PHB finished the job last week. TE has to confirm. We will not be able to assess whether this project was a success as of yet because the units are being operated 24/7 in that area due to other heating and cooling considerations.

01/24/05 – Peck started the job. They need a different transition and will be fabricating that in their shop and finishing the job in one or two days.

01/18/05 – Peck said they should start on Wed or Thur, Jan 19, 20 per Ken Shad.

01/03/05 – Peck should be onsite at the end of the week or beginning of next week to start after hours work.

12/24/04 – TE has given Peck a scheme to price up. Assuming they have the best price, they are the odds on favorite to do the work since they have an unused $5000 allowance at the CH in the Archive Room.

12/10/04 – Price came back from B&J at $6000, TE is now inviting Peck to give us a price since they have allowance money left on 2 County jobs. TTC will bring down Peck for a price quote next week. We will coordinate the visit with Mike Bartlett.

11/05/04 – TE met with Mike Bartlett, Bill Scholl & Jim (Bailiff’s) and reviewed the general layout of the 5-ton split system into the Bailiff’s area but the actual unit would be in the Room #167 space (this would have to be approved by LIP), ductwork would have to routed above the 167 space and over to the 169 area, 6 diffusers at 250 cfm would have to be installed to meet the 2000 cfm requirement of the 5-ton unit with a return ductwork. the final design is still under review. TE has asked B&J Jacobs to take a look at the design to provide their opinion about what is easiest and most cost effective. TTC will provide prices associated with each of these schemes so that the County can make an informed decision.

11/04/04 – TE discussed a ductless split system installation with Mike Bartlett, Bill Scholl, Diana Supe and left a message for Tony Matre in the 24 hour bailiff’s area today. TE will follow up with Tony about funding this project ASAP.

10/25/04 – Bill Scholl called today. A full size duct won’t fit through above the ceiling into the bailiff’s area due to a water pipe on the other side of the wall but a duct sized at 12″ x something will. I told them to find out and maximize the size of this duct and proceed with the design. We are trying to get as much air into the space as possible without spending much money. This is the first stage of the “fix”. Other things may have to be done later to enhance the space control.

10/14/04 – Mike Bartlett showed TTC B-12 which is actually the unit that serves the Concessions area. This unit is served by both heating hot water and chilled water. There is a temperature sensor in the supply duct outlet that monitors the leaving supply air temperature. TE believes that this is the sensor that Bill Scholl was talking about in our 10-12-04 meeting. When there is no building chilled water, the DX unit provides cooling for B-12. Our drawing and scheme will still work as the timeclock will divert air to the room #169 space starting at 7:00 p.m. – 6:00 a.m. and regardless of how B-12 is providing cool air it will get into 169 to assist in cooling the space. The system should still be modified as design on the ThermalTech packet.

10/12/04 – TTC met with DES, MB and Bill Scholl to discuss the reworking of the ductwork on unit B-9 to included Room #169 during the evening time. Trades and HVAC will do the work in the next couple of weeks. The plan will have B-9 provide cooling to the space after hours between 7pm and 6am. TE also stated that the ceiling work required should be done after hours as this will be a messy job especially in the kitchen area above the sink. We will monitor the temperature during the first few months and see if the County needs to add additional controls to the new air damper later. Bill reported that a Paragon timeclock will be used as they have one of those instock and the damper and actuator have been ordered. TE is fine with this information.

10/08/04 – TEI received a letter asking a question from the plan examiner. MS needs to talk with him because a letter will answer his question. However, he is out of the country till the 18th. They directed me to his supervisor Bill Spurling. TE have been playing phone tag with him for the past couple of days.

10/07/04 – TE visited the site and TTC has sketched up a possible solution to the Room #169 heating issue. TTC will deliver to the County and look at making the changes utilizing existing equipment for the work on 10-08-04.

09/13/04 – Fire Test at 6:00 a.m. Space did not pass. Egress into LIP has been questions by Larry Keifer. Security drawing was not included in permit set to City. Mike Stiles to rectify ASAP.

09/03/04 – TE delivered a set of permit prints to Brian Meyer for use on this project.

Security issues are ongoing between LIP and Recorders.

Trades doing all work on this projects.

10/18/04 – TEI received question from plan examiner and TEI responded. Plan examiner was out of the country until Monday October 18th. TE waiting for response.

4/19/05 – TEI recieved call from Mark Harden indicating that the State Agency that governs the computer system information will not accept the security level provided by the rear door in the computer room. However, the motion sensor that was added was per the the state building code. Mark and TEI to discuss options. TEI contacted Ben with GBBN to discuss the option of another 2nd path of egress. Ben to respond.